INAF queue observing, 2017 Jun 13-14

Observer: A. Rossi, F. Cusano
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC, LUCI

Summary

At the sunset the sky is clear.

Night duration (12\xB0 el twilight) 4:00am-10:30pm UT = 6 h 30m

Observing time: * *4.5hrs * *0m (70%)

Weather time loss: *0h *0m (0%)

technical time loss * 2.*hrs (30%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Bias   LBC     25 bino Bias, LBCred files have wrong name, in repository too
Sky Flat   LBC g,r   LBCred files have wrong name, in repository too
Annibali R34 UGC8320 LBC g,r 0.8h  
Fiorentino AO Palomar 10 LUCI AO J,Ks 1h completed
Antoniucci AO LKHa233 LUCI AO FeII, H2 0.5h  
Flats N30 camera LUCI Ks, J, H2, FeII    

Details

1:50 starting bino bias 25
  • NOTE we later found that LBCred fits files have date in filename: 20170114 (wrong month!),
    Also, they are in the repository 20170114

2:40 evening skyflats g, r for Annibali
  • NOTE we later found that LBCred fits files have date in filename: 20170114 (wrong month!),
    Also, they are in the repository 20170114

3:28 slewing to copoiting of Annibali R34; target UGC8320
  • 3:31 dohybrid, error no file found in new data
  • 3:35 dohybrid

3:30 start collimating dohybrid
  • 3:40 LBC red images with wrong month in filename. cannot proceed with dohybrid
  • 3:55 corrected date, running dohybrid

4:03 copointing Annibali

4:07 start Annibali R34, script UGC8320_1
  • 4:12 DIMM seeing 0.9, 1.0 on r image
  • 4:26 stars elongated in the blue channel we stop the ob at 6th exposure
  • 4:28 dofpia, collimation again
  • 4:30 send the script UGC8320_1 again
  • seeing 0.7" at 4:38

4:59 dofpia

5:04 started Annibali UGC8320_2
  • seeing 0.8" dimm
  • seeing 0.9" on images at 5:24

5:32 ob finished moon rise, we switch to LUCI

5:57 slewing to Rossi ESM program, GRB130803A
  • aborted the script, AO preset still not running

6:01 started again Rossi program, GRB130803A
  • 6:04 error during comitting setup, problem with the actuator
  • 6:14 send the script again
  • 6:19 failed acquisition again, Doug Miller is checking

6:38 we try to put the mask in FPU, because in turn-out it warms up LUCI, but we had a problem to commit the setup, the mask is in unknow state

7:25 mask problem solved

7:25 slew to Fiorentino LUCI AO program, Palomar 10 seeing 0.6 !!
  • luci2.20170614.0052-0193

8:17 Pal 10 seeing limited, Fiorentino
  • luci2.0194-0199 not good
  • luci2.0200-0209 guider probe in the field

8:47 starting again Fiorentio Palomar 10
  • the Arcetri group and Doug Miller applied a correction for astigmatism after checking the images
  • luci2.0210-0351
  • seeing 0.8" on the guider a 9:03

9:31 seeing limited Pal 10
  • luci2.0352-0361

9:44 checking the alignment of the N30 camera
  • 5min lost due to gui frozen

9:51 slewing to Antoniucci AO, LKHa233
  • 5min lost due to OT script naming problem (AO ref cannot be same as target)
  • luci2.0363-0384 pause to correct offsets (images are good)
  • seeing 0.4 on guider at 10:38 !
  • luci2.0385-0434 run again H2 starting with luci2.0385-0411
  • seeing 0.5 on guider at 11:08

11:20 slewing to PSF standard of Antoniucci
  • luci2.0435-0436 stop loop not closing at 11:32 too high background

11:35 END SCIENCE

* Skyflats, N30 camera
  • luci2.0437-0449 J band , first 5 have low counts
  • luci2.0460-0470 Ks band

* Dome flats, N30 camera
  • luci2.0471-0475 Ks too low
  • luci2.0482- Ks good
  • luci2.0487-0496 Ks good
  • luci2.0499-0507 H2 good
  • luci2.0510-0519 FeII good
  • luci2.0520-0529 J good (johny be goode)

*dome flats, camera, N3.75
  • luci2.0531-0540 J good
  • luci2.0542-0551 Ks good

-- %USERSIG{AndreaRossi - 2017-06-14}%

Comments

 
Topic revision: r19 - 09 Nov 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback